OpenStudy (anonymous):

there is a useful grass and weed cutting tool that utilizes a 0.21-m length of nylon "string" rotating at 6200 rev/min about an axis perpendicular to one end of the string. (a) What is the time needed for the string to sweep out an angle of 35 degree? (b) Assuming that the length of the string does not change, find the distance through which the tip of the string moves during this interval.

OpenStudy (mos1635):

6200 rev/min = frequency of rotation do you know formula for angular velocity ω ?

OpenStudy (anonymous):

ω = θ / t

OpenStudy (mos1635):

wright! but that is the definition of ω you will use it after you have calculate the ω now formula for ω involving frequency???

OpenStudy (anonymous):

I don't know.

OpenStudy (mos1635):

ω=2 π f

OpenStudy (mos1635):

note you have to convert rev/min to rev/sec

OpenStudy (anonymous):

Alright.

OpenStudy (mos1635):

so we have ω=2*π* f ω=θ/t there for 2*π*f=θ/t t=θ/2πf f to rev/sec and θ to rad

OpenStudy (anonymous):

Thanks let's see if I can solve it now.

OpenStudy (mos1635):

you should find t=9.4086exp-4 sec

OpenStudy (anonymous):

This is what I got for the first equation 6200 rev/min = 103.3 rev/s ω = 2π(103.3) ω=648.724

OpenStudy (anonymous):

Am I correct ?

OpenStudy (mos1635):

what you use for π ?

OpenStudy (anonymous):

pi= 3.14

OpenStudy (mos1635):

ok you are correct

OpenStudy (mos1635):

how ever you could have a more acurate resalt if t=θ/2πf θ=(35/360)*2π rad f=6200/60 Hz t=....... t=(35*2π*60)/360*2π*6200 ...t=9.4086exp-4 sec this way you use your calculator only once at the end and have only one aproximation

OpenStudy (mos1635):

in my place of the planet that is the way we do calculations don;t know if that is axeptable for you...

OpenStudy (anonymous):

My teacher didn't teach me that way

OpenStudy (mos1635):

ok folow you way This is what I got for the first equation 6200 rev/min = 103.3 rev/s ω = 2π(103.3) ω=648.724 rad/sec so far OK

OpenStudy (mos1635):

now you have to convert degrees to rad do you know how?

OpenStudy (anonymous):

35 degree = 0.6108652381965 rad

OpenStudy (mos1635):

0.61055555555555555555rad

OpenStudy (mos1635):

in your calculator: 35 / 360 = * 2 * 3.14 = 0.61055555555

OpenStudy (mos1635):

now you have θ=0.610555 rad ω=648.724 rad/sec go for t

OpenStudy (anonymous):

Alright thanks

OpenStudy (anonymous):

I got 1062.5

OpenStudy (mos1635):

for your next question